# Convergence order of the quantization error for self-affine measures on Lalley-Gatzouras carpets

### v1

@@ 2.2. Characterizations for the quantization error
(i) First we assume that $\sigma_{L}=\omega_{L}$, then $\sigma_{R},\omega_{R}$ are incomparable. Note that
$\overline{\sigma}_{y}|_{k_{1}+2}=(\sigma_{L})_{y}\ast(j_{0},j_{0})\ast\sigma_{R},\;\overline{\omega}_{y}|_{k_{2}+2}=(\sigma_{L})_{y}\ast(j_{0},j_{0})\ast\omega_{R}.$
It follows that $\overline{\sigma}_{y},\overline{\omega}_{y}$ are incomparable.
@@
$\overline{\omega}:=\sigma_{L}\ast(\hat{i}_{l+1},\hat{j}_{l+1})\ast\tau_{0}\ast(\hat{j}_{l+2},\ldots,\hat{j}_{k_{2}},\ldots,j_{\overline{k}_{2}}).$
@@
$\overline{\omega}_{y}=(j_{1},\ldots,j_{l},\hat{j}_{l+1},j_{0},j_{0},\hat{j}_{l+2},\ldots,\hat{j}_{k_{2}},\ldots,j_{\overline{k}_{2}}).$
If $\hat{j}_{l+1}\neq j_{0}$, then $\overline{\sigma}_{y},\overline{\omega}_{y}$ are incomparable. Next, we assume that $\hat{j}_{l+1}=j_{0}$. If $j_{l+1}\neq j_{0}$, we again have that $\overline{\sigma}_{y},\overline{\omega}_{y}$ are incomparable; otherwise, we have, $\hat{j}_{l+1}=j_{l+1}=j_{0}$. Note that $\sigma_{y},\omega_{y}$ are incomparable. We deduce that $(j_{l+2},\ldots,j_{k_{1}})$ and $(\hat{j}_{l+2},\ldots,\hat{j}_{k_{2}})$ are incomparable. Hence, $\overline{\sigma}_{y},\overline{\omega}_{y}$ are incomparable.
(iii) Finally, we assume that $|\omega_{L}|=|\sigma_{L}|+2$. We write
$\overline{\sigma}:=\sigma_{L}\ast\tau_{0}\ast(j_{l+1},j_{l+2}\ldots,j_{k_{1}},\ldots,j_{\overline{k}_{1}});$
$\overline{\omega}:=\sigma_{L}\ast((\hat{i}_{l+1},\hat{j}_{l+1}),(\hat{i}_{l+2},\hat{j}_{l+2}))\ast\tau_{0}\ast(\hat{j}_{l+3},\ldots,\hat{j}_{k_{2}},\ldots,j_{\overline{k}_{2}}).$
Then $\overline{\sigma}_{y}$ and $\overline{\omega}_{y}$ take the following form:
$\overline{\sigma}_{y}=(j_{1},\ldots,j_{l},j_{0},j_{0},j_{l+1},j_{l+2},j_{l+3},\ldots,j_{k_{1}},\ldots,j_{\overline{k}_{1}}),$
$\overline{\omega}_{y}=(j_{1},\ldots,j_{l},\hat{j}_{l+1},\hat{j}_{l+2},j_{0},j_{0},\hat{j}_{l+3},\ldots,\hat{j}_{k_{2}},\ldots,j_{\overline{k}_{2}}).$
If $(\hat{j}_{l+1},\hat{j}_{l+2})\neq(j_{0},j_{0})$, then $\overline{\sigma}_{y},\overline{\omega}_{y}$ are incomparable. In the following, we assume that $(\hat{j}_{l+1},\hat{j}_{l+2})=(j_{0},j_{0})$. If $(j_{l+1},j_{l+2})\neq(j_{0},j_{0})$, then $\overline{\sigma}_{y},\overline{\omega}_{y}$ are incomparable. Finally, we assume that $(\hat{j}_{l+1},\hat{j}_{l+2})=(j_{l+1},j_{l+2})=(j_{0},j_{0})$. Because $\sigma_{y},\omega_{y}$ are incomparable, we deduce that $(j_{l+3},\ldots,j_{k_{1}})$ and $(\hat{j}_{l+3},\ldots,\hat{j}_{k_{2}})$ are incomparable. This implies that $\overline{\sigma}_{y},\overline{\omega}_{y}$ are again incomparable, hence the lemma follows. ∎
As the following example shows, things are getting more complicated, when $\sigma_{L},\omega_{L}$ are comparable and $\big{|}|\sigma_{L}|-|\omega_{L}|\big{|}\geq 3$.
Example 2.13.
We assume that $\sigma_{L}\precneqq\omega_{L}$ and $|\omega_{L}|=|\sigma_{L}|+3$. At this moment we temporally do not require that $\sigma,\omega\in\Lambda_{n,r}$. Next, we show that, it is possible that $\sigma_{y},\omega_{y}$ are incomparable, but $\overline{\sigma}_{y},\overline{\omega}_{y}$ are comparable. We write
$\sigma=\sigma_{L}\ast(j_{l+1},j_{0},j_{0},j_{l+4},j_{l+5},\ldots,j_{k_{1}});$
$\omega=\sigma_{L}\ast((\hat{i}_{l+1},j_{0}),(\hat{i}_{l+2},j_{0}),(\hat{i}_{l+3},\hat{j}_{l+3}))\ast(\hat{j}_{l+4},\ldots,\hat{j}_{k_{2}}).$
Here, we have assumed that $j_{l+2}=j_{l+3}=j_{0}=\hat{j}_{l+1}=\hat{j}_{l+2}$. We have
$\sigma_{y}=(j_{1},\ldots,j_{l},j_{l+1},j_{0},j_{0},j_{l+4},\ldots,j_{k_{1}}),$
$\omega_{y}=(j_{1},\ldots,j_{l},j_{0},j_{0},\hat{j}_{l+3},\hat{j}_{l+4},\ldots,\hat{j}_{k_{2}}).$
When $j_{l+1}\neq j_{0}$, one can see that $\sigma_{y},\omega_{y}$ are incomparable, but
$\overline{\sigma}_{y}=(j_{1},\ldots,j_{l},j_{0},j_{0},j_{l+1},j_{0},j_{0},j_{l+4},\ldots,j_{k_{1}},\ldots,j_{\overline{k}_{1}}),$
$\overline{\omega}_{y}=(j_{1},\ldots,j_{l},j_{0},j_{0},\hat{j}_{l+3},j_{0},j_{0},\hat{j}_{l+4},\ldots,\hat{j}_{k_{2}},\ldots,j_{\overline{k}_{2}}).$
If $\hat{j}_{l+3}=j_{l+1}$, and the two words $(j_{l+4},\ldots,j_{k_{1}},\ldots,j_{\overline{k}_{1}}),(\hat{j}_{l+4},\ldots,\hat{j}_{k_{2}},\ldots,j_{\overline{k}_{2}})$ are comparable, then $\overline{\sigma}_{y},\overline{\omega}_{y}$ are comparable.
For the words $\sigma,\omega$ in Example 2.13, we will show in the next lemma that $k_{1}\leq k_{2}$. It is helpful to note that $\omega_{y}|_{k_{1}}$ is a permutation of $\sigma_{y}$. This will actually cross out the possibility that $\sigma,\omega$ belong to $\Lambda_{n,r}$ simultaneously, when $n$ is sufficiently large. Next, we will prove by contradiction that, for sufficiently large $n$ and for every pair of words $\sigma,\omega\in\Lambda_{n,r}$ with $\sigma_{L},\omega_{L}$ comparable, $\overline{\sigma}_{y},\overline{\omega}_{y}$ are necessarily incomparable.
@@
Let $A_{5,r}:=\big{[}\frac{\log(\underline{q}^{2}\underline{\eta}_{r})}{r\log\overline{a}}\big{]}$ and $n\geq T_{1,r}:=[A_{4}^{-1}(A_{1}+A_{5,r}+3)]+1$. Then for $\sigma,\omega\in\Lambda_{n,r}$ with $\sigma_{L}\precneqq\omega_{L}$, we have that $\overline{\sigma}_{y},\overline{\omega}_{y}$ are incomparable.
@@
Assume that $\sigma,\omega\in\Lambda_{n,r},\sigma_{L}\precneqq\omega_{L}$, but $\overline{\sigma}_{y},\overline{\omega}_{y}$ are comparable. We will deduce a contradiction. Let $\sigma,\omega$ be the same as in (2.4) and (2.5) with $h\geq 3$ (cf. Lemma 2.12). Then
$\overline{\sigma}:=\sigma_{L}\ast\tau_{0}\ast(j_{l+1},\ldots,j_{k_{1}},\ldots,j_{\overline{k}_{1}});$
$\overline{\omega}:=\sigma_{L}\ast((\hat{i}_{l+1},\hat{j}_{l+1}),\hat{i}_{l+2},\hat{j}_{l+2})\ldots,(\hat{i}_{l+h},\hat{j}_{l+h}))\ast\tau_{0}\ast(\hat{j}_{l+h+1},\ldots,\hat{j}_{k_{2}},\ldots,j_{\overline{k}_{2}}).$
@@
$\overline{\omega}_{y}=(j_{1},\ldots,j_{l},\hat{j}_{l+1},\hat{j}_{l+2},\hat{j}_{l+3},\hat{j}_{l+4},\ldots,\hat{j}_{l+h},j_{0},j_{0},\hat{j}_{l+h+1},\ldots,\hat{j}_{k_{2}},\ldots,j_{\overline{k}_{2}}).$
@@
Note that $\sigma\in\Psi_{l}$. we have $\prod_{h=1}^{k_{1}-1}b_{j_{h}}\geq a_{\sigma_{L}}>\prod_{h=1}^{k_{1}}b_{j_{h}}$. Hence,
$b_{\omega_{y}}$ $=$ $\prod_{h=1}^{l}b_{j_{h}}\prod_{h=l+1}^{l+h}b_{\hat{j}_{h}}\prod_{h=l+h+1}^{k_{2}}b_{\hat{j}_{h}}$
$=$ $\prod_{h=1}^{l}b_{j_{h}}\bigg{(}b_{j_{0}}^{2}\prod_{h=l+3}^{l+h}b_{\hat{j}_{h}}\bigg{)}\prod_{h=l+h+1}^{k_{2}}b_{\hat{j}_{h}}$
$=$ $\prod_{h=1}^{l}b_{j_{h}}\bigg{(}\prod_{h=l+1}^{l+h-2}b_{j_{h}}b_{j_{0}}^{2}\bigg{)}\prod_{h=l+h+1}^{k_{2}}b_{j_{h}}$
$\geq$ $\prod_{h=1}^{k_{1}-1}b_{j_{h}}>a_{\sigma_{L}}>a_{\omega_{L}}.$
@@
(2.8) $(\hat{j}_{l+h+1},\ldots,\hat{j}_{k_{1}})\preceq(j_{l+h+1},\ldots,j_{k_{2}})\;\;{\rm if}\;k_{1}>l+h.$
@@
$\mathcal{E}_{r}(\omega)$ $=$ $\big{(}p_{\sigma_{L}}\prod_{h=l+1}^{l+h}p_{\hat{i}_{h}\hat{j}_{h}}\prod_{h=l+h+1}^{k_{2}}q_{\hat{j}_{h}}\big{)}(a_{\sigma_{L}}^{r}\prod_{h=l+1}^{l+h}a_{\hat{i}_{h}\hat{j}_{h}}^{r})$
$=$ $\big{(}p_{\sigma_{L}}p_{\hat{i}_{l+1}\hat{j}_{l+1}}p_{\hat{i}_{l+2}\hat{j}_{l+2}}\prod_{h=l+1}^{l+h-2}p_{i_{h}j_{h}}\prod_{h=l+h+1}^{k_{2}}q_{\hat{j}_{h}}\big{)}(a_{\sigma_{L}}^{r}\prod_{h=l+1}^{l+h}a_{\hat{i}_{h}\hat{j}_{h}}^{r})$
$\leq$ $\big{(}p_{\sigma_{L}}\prod_{h=l+1}^{l+h-2}q_{j_{h}}\prod_{h=l+h+1}^{k_{1}}q_{j_{h}}\big{)}(a_{\sigma_{L}}^{r}\prod_{h=l+1}^{l+h}a_{\hat{i}_{h}\hat{j}_{h}}^{r})$
@@
Case 2: $k_{1}\leq l+h$. In this case, we have
$\mathcal{E}_{r}(\omega)$ $\leq$ $\big{(}p_{\sigma_{L}}\prod_{h=l+1}^{l+h-2}q_{j_{h}}\big{)}(a_{\sigma_{L}}^{r}\prod_{h=l+1}^{l+h}a_{\hat{i}_{h}\hat{j}_{h}}^{r})$
$\leq$ $\big{(}p_{\sigma_{L}}\prod_{h=l+1}^{k_{1}-2}q_{j_{h}}\big{)}(a_{\sigma_{L}}^{r}\prod_{h=l+1}^{l+h}a_{\hat{i}_{h}\hat{j}_{h}}^{r})$
@@
Using Claims 1-3, we are able to complete the proof of the lemma. We again distinguish between two cases.
Case (i): $A_{5,r}<3$. In this case, we have, $3\leq h\leq A_{5,r}<3$, a contradiction.
Case (ii): $A_{5,r}\geq h\geq 3$. Because $n\geq T_{1,r}$, by Lemma 2.6 and Remark 2.7 (2), we have, $|\sigma_{R}|=k_{1}-l\geq A_{1}+A_{5,r}+3\geq A_{1}+h+3$. Thus,
@@
We define $\breve{\omega}:=\sigma_{L}\ast(j_{0},j_{0},\hat{j}_{l+3},\ldots,\hat{j}_{l+h},\hat{j}_{l+h+1},\ldots,\hat{j}_{k_{1}})$. Then we have $F_{\omega}\subset F_{\breve{\omega}}$. Since $k_{1}>l+h$, by (2.6)-(2.8), $(\breve{\omega})_{R}$ is a permutation of $\sigma_{R}$. Note that, by Lemma 2.4 (ii), we have, $|\sigma^{\flat}|\geq|\sigma|-A_{1}\geq A_{5,r}+3>h+2$. Thus, $(\breve{\omega}^{\flat})_{R}$ is also a permutation of $(\sigma^{\flat})_{R}$. It follows that
@@
In the remaining part of this section, we assume that $n\geq T_{2,r}:=T_{1,r}+2A_{2}$. Let $\overline{\sigma}\in\mathcal{B}_{n,r}$ be as defined in (2.3). For every $1\leq h\leq 2A_{2}$ , we fix an integer $i_{l+1+h}\in[1,n_{j_{l+1+h}}]$ and define (cf. Lemma 2.4 (ii))
@@
Then we have $F_{\overline{\sigma}^{*}}\subset F_{\overline{\sigma}}$. We define
(2.9) $\mathcal{F}_{n,r}:=\{\overline{\sigma}*:\overline{\sigma}\in\mathcal{B}_{n,r}\}.$
Remark 2.15.
(1) From the definitions of $\mathcal{B}_{n,r}$ and $\mathcal{F}_{n,r}$, one can easily see that, ${\rm card}(\mathcal{F}_{n,r})={\rm card}(\mathcal{B}_{n,r})=\varphi_{n,r}$. (2) Let $K_{n,r}:=\bigcup_{\overline{\sigma}^{*}\in\mathcal{F}_{n,r}}F_{\overline{\sigma}^{*}}$. By the definition of $\overline{\sigma}^{*}$, we have, $|\overline{\sigma}^{*}_{L}|-|\sigma_{L}|\leq 2+2A_{2}$. This and Lemma 2.4 (i) yield that
$\mu(K_{n,r})\geq\underline{p}^{2(A_{2}+1)(1+A_{1})}\sum_{\sigma\in\Lambda_{n,r}}\mu(F_{\sigma})\geq\underline{p}^{8A_{1}A_{2}}.$
(3) As we showed in [20, Lemma 4], there exists a positive number $D_{L}$, such that, for every $\alpha_{L}\subset\mathbb{R}^{2}$ with cardinality $L$, the following holds:
@@
$d(F_{\overline{\sigma}^{*}},F_{\overline{\omega}^{*}})\geq(1+\underline{b}^{-2})^{-1}\underline{b}^{2}\max\{|F_{\overline{\sigma}^{*}}|,|F_{\overline{\omega}^{*}}|\}.$
@@
If $(\overline{\sigma})_{y},(\overline{\omega})_{y}$ are comparable, then by Lemmas 2.12, 2.14, $\sigma_{L},\omega_{L}$ are incomparable. Thus, from Lemma 2.11, we have
@@
With the above preparations, we can now establish a characterization for the quantization error by applying [20, Lemma 3].
@@
By (2.10), Lemma 2.16 and Remark 2.15 (3), the assumptions in Lemma 3 of [20] are fulfilled for the conditional measure $\mu_{n,r}:=\mu(\cdot|K_{n,r})$. It follows that

### v10

@@ 2.2. Characterizations for the quantization error
(i) First we assume that $\sigma_{L}=\omega_{L}$, then $\sigma_{R},\omega_{R}$ are incomparable. Note that $\overline{\sigma}_{y}|_{k_{1}+2}=(\sigma_{L})_{y}\ast(j_{0},j_{0})\ast\sigma_{R}$ and $\overline{\omega}_{y}|_{k_{2}+2}=(\sigma_{L})_{y}\ast(j_{0},j_{0})\ast\omega_{R}$. It follows that $\overline{\sigma}_{y},\overline{\omega}_{y}$ are incomparable.
@@
$\overline{\omega}:=\sigma_{L}\ast(\hat{i}_{l+1},\hat{j}_{l+1})\ast\tau_{0}\ast(\hat{j}_{l+2},\ldots,\hat{j}_{k_{2}},\ldots,\hat{j}_{\overline{k}_{2}}).$
@@
$\overline{\omega}_{y}=(j_{1},\ldots,j_{l},\hat{j}_{l+1},j_{0},j_{0},\hat{j}_{l+2},\ldots,\hat{j}_{k_{2}},\ldots,\hat{j}_{\overline{k}_{2}}).$
If $\hat{j}_{l+1}\neq j_{0}$, then $\overline{\sigma}_{y},\overline{\omega}_{y}$ are incomparable. Next, we assume that $\hat{j}_{l+1}=j_{0}$. If $j_{l+1}\neq j_{0}$, we again have that $\overline{\sigma}_{y},\overline{\omega}_{y}$ are incomparable; otherwise, we have, $\hat{j}_{l+1}=j_{l+1}=j_{0}$. Note that $\sigma_{y},\omega_{y}$ are incomparable. Thus, $(j_{l+2},\ldots,j_{k_{1}})$ and $(\hat{j}_{l+2},\ldots,\hat{j}_{k_{2}})$ are incomparable. Hence, $\overline{\sigma}_{y},\overline{\omega}_{y}$ are incomparable.
(iii) The case that $|\omega_{L}|=|\sigma_{L}|+2$ can be proved analogously to (ii). ∎
Next, we prove by contradiction that, for $\sigma,\omega\in\Lambda_{n,r}$ with $\sigma_{L}\precneqq\omega_{L}$ and $|\omega_{L}|-|\sigma_{L}|\geq 3$, $\overline{\sigma}_{y},\overline{\omega}_{y}$ are incomparable: assuming the contrary, we will construct some word $\breve{\omega}\precneqq\omega$ such that $\breve{\omega}\in\Lambda_{n,r}$, contradicting the fact that $\Lambda_{n,r}$ is a finite anti-chain.
@@
Let $A_{5,r}:=\big[\frac{\log(\underline{q}^{2}\underline{\eta}_{r})}{r\log\overline{a}}\big]$ and $n>T_{1,r}:=[A_{4}^{-1}(A_{1}+A_{5,r}+3)]$. Let $\sigma,\omega\in\Lambda_{n,r}$. Assume that $\sigma_{L}\precneqq\omega_{L}$ and $|\omega_{L}|-|\sigma_{L}|\geq 3$. Then $\overline{\sigma}_{y},\overline{\omega}_{y}$ are incomparable.
@@
Assume that $\sigma,\omega\in\Lambda_{n,r},\sigma_{L}\precneqq\omega_{L}$, but $\overline{\sigma}_{y},\overline{\omega}_{y}$ are comparable. We will deduce a contradiction. Let $\sigma,\omega$ be the same as in (2.7) and (2.8) with $h\geq 3$. Then we have $\overline{\sigma}:=\sigma_{L}\ast\tau_{0}\ast(j_{l+1},\ldots,j_{k_{1}},\ldots,j_{\overline{k}_{1}})$ and
$\overline{\omega}:=\sigma_{L}\ast((\hat{i}_{l+1},\hat{j}_{l+1}),\hat{i}_{l+2},\hat{j}_{l+2})\ldots,(\hat{i}_{l+h},\hat{j}_{l+h}))\ast\tau_{0}\ast(\hat{j}_{l+h+1},\ldots,\hat{j}_{k_{2}},\ldots,\hat{j}_{\overline{k}_{2}}).$
@@
$\overline{\omega}_{y}=(j_{1},\ldots,j_{l},\hat{j}_{l+1},\hat{j}_{l+2},\hat{j}_{l+3},\hat{j}_{l+4},\ldots,\hat{j}_{l+h},j_{0},j_{0},\hat{j}_{l+h+1},\ldots,\hat{j}_{k_{2}},\ldots,\hat{j}_{\overline{k}_{2}}).$
@@
Note that $\sigma\in\Psi_{l}$. we have $\prod_{p=1}^{k_{1}-1}b_{j_{p}}\geq a_{\sigma_{L}}>\prod_{p=1}^{k_{1}}b_{j_{p}}$. Hence,
$b_{\omega_{y}}$ $=$ $\prod_{p=1}^{l}b_{j_{p}}\prod_{p=l+1}^{l+h}b_{\hat{j}_{p}}\prod_{p=l+h+1}^{k_{2}}b_{\hat{j}_{p}}$
[…]
$\geq$ $\prod_{p=1}^{k_{1}-1}b_{j_{p}}\geq a_{\sigma_{L}}>a_{\omega_{L}}.$
@@
(2.11) $(j_{l+h+1},\ldots,j_{k_{1}})\preceq(\hat{j}_{l+h+1},\ldots,\hat{j}_{k_{2}})\;\;{\rm if}\;k_{1}>l+h.$
@@
$\mathcal{E}_{r}(\omega)$ $=$ $\big(p_{\sigma_{L}}\prod_{p=l+1}^{l+h}p_{\hat{i}_{p}\hat{j}_{p}}\prod_{p=l+h+1}^{k_{2}}q_{\hat{j}_{p}}\big)(a_{\sigma_{L}}^{r}\prod_{p=l+1}^{l+h}a_{\hat{i}_{p}\hat{j}_{p}}^{r})$
$\leq$ $\big(p_{\sigma_{L}}p_{\hat{i}_{l+1}\hat{j}_{l+1}}p_{\hat{i}_{l+2}\hat{j}_{l+2}}\prod_{p=l+3}^{l+h}q_{\hat{j}_{p}}\prod_{p=l+h+1}^{k_{2}}q_{\hat{j}_{p}}\big)(a_{\sigma_{L}}^{r}\prod_{p=l+1}^{l+h}a_{\hat{i}_{p}\hat{j}_{p}}^{r})$
[…]
@@
Case 2: $k_{1}\leq l+h$. If $k_{1}\geq l+h-2$, then 2.10) remains valid; otherwise, we have $(j_{l+1},\ldots,j_{k_{1}})=(\hat{j}_{l+3},\ldots,\hat{j}_{k_{1}+3})$. Let $k:=\min(k_{1},l+h-2)$, we deduce
$\mathcal{E}_{r}(\omega)$ $\leq$ $\big(p_{\sigma_{L}}\prod_{p=l+1}^{k}q_{j_{p}}\big)(a_{\sigma_{L}}^{r}\prod_{p=l+1}^{l+h}a_{\hat{i}_{p}\hat{j}_{p}}^{r})$
$\leq$ $\big(p_{\sigma_{L}}\prod_{p=l+1}^{k_{1}-2}q_{j_{p}}\big)(a_{\sigma_{L}}^{r}\prod_{p=l+1}^{l+h}a_{\hat{i}_{p}\hat{j}_{p}}^{r})$
@@
Using Claims 1-3, we are able to complete the proof of the lemma. Because $n>T_{1,r}$, by Lemma 2.3 and Remark 2.4 (2), we have,
(2.12) $|\sigma_{R}|=k_{1}-l\geq A_{1}+A_{5,r}+3\geq A_{1}+h+3.$
Thus, we may write $\sigma_{y},\omega_{y}$ in detail:
@@
Let $\breve{\omega}:=\sigma_{L}\ast(j_{0},j_{0},\hat{j}_{l+3},\ldots,\hat{j}_{l+h},\hat{j}_{l+h+1},\ldots,\hat{j}_{k_{1}})$. By (2.9)-2.12, $(\breve{\omega})_{R}\;({\rm respectively}\;(\breve{\omega})_{R}^{-})$ is a permutation of $\sigma_{R}\;({\rm respectively}\;\sigma_{R}^{-})$. Thus,
@@
It follows that $\breve{\omega}\in\Psi_{l}$ and $F_{\omega}\subsetneq F_{\breve{\omega}}$. Note that, by Lemma 2.2 (ii), we have, $|(\sigma^{\flat})_{R}|\geq|\sigma_{R}|-A_{1}\geq A_{5,r}+3\geq h+3$. For some integer $T\geq 2$, we may write
$(\sigma^{\flat})_{R}=(j_{l},j_{l+1},\ldots,j_{l+h-2},j_{0},j_{0},j_{l+h+1},\ldots,j_{l+h+T}).$
Again, by (2.9)-(2.11) and the definition of $\Psi_{l-1}$, we deduce
$(\breve{\omega}^{\flat})_{R}=(j_{l},j_{0},j_{0},\hat{j}_{l+3},\ldots,\hat{j}_{l+h},\hat{j}_{l+h+1},\ldots,\hat{j}_{l+h+T}).$
Thus, $(\breve{\omega}^{\flat})_{R}$ is also a permutation of $(\sigma^{\flat})_{R}$. It follows that
@@
In the remaining part of this section, we assume that $n>T_{2,r}:=T_{1,r}+2A_{4}^{-1}A_{2}$. Let $\overline{\sigma}\in\mathcal{B}_{n,r}$ be as defined in (2.6). For every $1\leq h\leq 2A_{2}$ , we fix an integer $i_{l+h}\in[1,n_{j_{l+h}}]$ and define (cf. Lemma 2.2)
$\overline{\sigma}^{*}_{L}:=\sigma_{L}\ast\tau_{0}\ast\big((i_{l+1},j_{l+1}),\ldots,(i_{l+2A_{2}},j_{l+2A_{2}})\big),$
@@
Let $\overline{\sigma}^{*}:=\overline{\sigma}^{*}_{L}\ast\overline{\sigma}^{*}_{R}$. Then we have $F_{\overline{\sigma}^{*}}\subset F_{\overline{\sigma}}$. We define
(2.13) $\mathcal{F}_{n,r}:=\{\overline{\sigma}^{*}:\overline{\sigma}\in\mathcal{B}_{n,r}\}.$
@@
$d(F_{\overline{\sigma}^{*}},F_{\overline{\omega}^{*}})\geq(1+\underline{b}^{-2})^{-1}\underline{a}^{2}\max\{|F_{\overline{\sigma}^{*}}|,|F_{\overline{\omega}^{*}}|\}.$
@@
If $(\overline{\sigma})_{y},(\overline{\omega})_{y}$ are comparable, then by Lemmas 2.9, 2.10, $\sigma_{L},\omega_{L}$ are incomparable. Thus, from Lemma 2.8, we have
@@
Remark 2.12.
(1) From the definitions of $\mathcal{B}_{n,r}$ and $\mathcal{F}_{n,r}$, one can easily see that, ${\rm card}(\mathcal{F}_{n,r})={\rm card}(\mathcal{B}_{n,r})=\varphi_{n,r}$. (2) Let $K_{n,r}:=\bigcup_{\overline{\sigma}^{*}\in\mathcal{F}_{n,r}}F_{\overline{\sigma}^{*}}$. By the definition of $\overline{\sigma}^{*}$, $|\overline{\sigma}^{*}_{L}|-|\sigma_{L}|=2+2A_{2}$. Thus, by Lemmas 2.2 and 2.11,
$\mu(K_{n,r})=\sum_{\sigma\in\Lambda_{n,r}}\mu(F_{\overline{\sigma}^{*}})\geq\underline{p}^{2(A_{2}+1)(1+A_{1})}\sum_{\sigma\in\Lambda_{n,r}}\mu(F_{\sigma})\geq\underline{p}^{8A_{1}A_{2}}.$
(3) As we showed in [13, Lemma 4], there exists a positive number $D_{L}$, such that, for every $\alpha_{L}\subset\mathbb{R}^{2}$ with cardinality $L$, the following holds:
With the above preparations, we can now establish a characterization for the quantization error by applying [13, Lemma 3].
@@
By (2.14), Lemma 2.11 and Remark 2.12 (3), the assumptions in Lemma 3 of [13] are fulfilled for the measure $\mu_{n,r}:=\mu(\cdot|K_{n,r})$. It follows that